OpenStudy (anonymous):

Joy has a piece of ribbon that is 6 7/8 in. long. She purchased a piece that is 35 7/8 in. long. Which expression should Joy use to estimate how many total inches of ribbon she has altogether? A. 36-7 B. 36 1/2-7 1/2 C. 36+7 D. 36 1/2+7 1/2

OpenStudy (anonymous):

no not a d.

OpenStudy (anonymous):

\[6 \frac{ 7 }{ 8 }+35\frac{ 7 }{ 8 }=6+35+\frac{ 7 }{ 8 }+\frac{ 7 }{ 8 }=41+\frac{ 7+7 }{ 8 }=41+\frac{ 7 }{ 4 }=41+1+\frac{ 3 }{ 4 }\] \[=42\frac{ 3 }{ 4 }\] so i think c is near.

OpenStudy (anonymous):

ok.

OpenStudy (anonymous):

can you help me with 1 more?

OpenStudy (anonymous):

A trail is 3 2/10 km long. Cassidy has run 9/10 of the distance of the trail. About how far has Cassidy run? Estimate by first rounding each number to the nearest 1/2 . A. 2km B. 3km C. 4km D. 7km

OpenStudy (anonymous):

b.?

OpenStudy (anonymous):

ik i have to do this Rounding to half is done as follows: 1. multiply the number by two. 2. round to the nearest integer. 3. Divide the result by two. Example: round 5 1/3 to the nearest half 1. 5 1/3 *2 = 10 2/3 2. nearest integer is 11 3. divide by two to get 5 1/2

OpenStudy (anonymous):

wait is it B or c? becasue 3.5 = 3 1/2 and it is closer to - 4

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

\[3+\frac{ 2 }{ 10 }=3+\frac{ 1 }{ 5 }=3+0=3(rounded \to 1/2)\] \[\frac{ 9 }{ 10 }=\frac{ 5 }{ 10 }+\frac{ 4 }{ 10 }=\frac{ 1 }{ 2 }+\frac{ 1 }{ 2 },(\frac{ 1 }{ 2 }*\frac{ 5 }{ 5 } =\frac{ 5 }{10 },near \frac{ 4 }{ 10 })\] =1 1 of 3=3km

OpenStudy (anonymous):

:o

OpenStudy (anonymous):

ty.

OpenStudy (anonymous):

let me see what my grade it

OpenStudy (anonymous):

100% :D ty my teacher just graded it and said that i did really good on my home work.

OpenStudy (anonymous):

i normaly do get 100's but still. i was confused with this ty.

OpenStudy (anonymous):

fan and metal for 100%